Latesha​, a single​ taxpayer, had the following income and deductions for the tax year 2021.

Compute Latesha's taxable income and federal tax liability for 2021 ​(round to dollars and ignore the qualified business income deduction and​ self-employment taxes for this​ problem).

Salary $100,000

Business income $25,000

Interest income from taxable bonds $10,000

Interest income from tax-exempt bonds $5,000

Total income from whatever source derived $140,000

Minus: Exclusions, as provided in the tax law

 Interest income from tax-exempt bonds -($5,000)

Gross income $135,000

Minus: Deductions for Adjusted gross income

 Business expenses $?

Adjusted gross income (AGI) $?

Minus: Deductions from AGI:

 Account name? $?

Account name? $?

Taxable income $?

Tax $?


1
Expert's answer
2022-01-30T14:11:39-0500


Latesha Taxable income = "Salary+Interest from taxable bonds"

Salary = $ 100,000

Interest from taxable bonds = $ 10,000

"Taxable income = 100000+10000"

Taxable income = $ 100,000


Federal tax liability:

Assuming the below tax schedule;

"[0 -9,950]= 10\%"

"[9,951 - 40,525] = 12\%" 

"[40,526 - 86,375] = 22\%"

"[86,376 -164,925] = 24\%"

(Tina, 2022)

getting the tax liability;

"9,950\times .1=995"

"(40525-9950)\times.12=3,669"

"(86375-40525)\times.22=10,087"

"(100000-86375)\times.24=3,270"

Tax liability = "995+3669+10087+3270"

Tax liability = $ 18,021